OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2. y - 5 = 3/2(x + 1) y + 1 = 3/2(x - 5) y + 5 = -3/2(x - 1) y - 1 = 3/2(x + 5)

OpenStudy (ash2326):

Line passing through point (x1, y1) and slope m has equation given as \[\frac{y-y1}{x-x1}=m\] or \[y-y1=m(x-x1)\] here (x1, y1)= (-5, 1) and slope m= \(\large \frac 32\)

OpenStudy (ash2326):

@JackJones Can you try now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

is the answer d

OpenStudy (anonymous):

the last one

OpenStudy (ash2326):

Yeah, good work :D

OpenStudy (anonymous):

thanks for giving me the formula